OpenStudy (anonymous):

I Have the function (ln(x)-1)^(1/2) determine the range and inverse of the function. I determined the inverse to be: e^(x^(2)-1) How do I find the domain and range of this function can someone explain this to me I'm lost

OpenStudy (watchmath):

why you deleted your post before. Your domain was correct. But the range is \([0,\infty)\). Since \(\ln(x)-1\) take values from \(-\infinity\) to \(\infinity\). So square root of that (the one that make sense) can take value from 0 to infinity.

OpenStudy (anonymous):

You can't input 0 or any negative value into ln(x) it is not possible

OpenStudy (anonymous):

I came up with the domain of [e, +infinity) and a Range of [(2)^(1/2), infinity)

OpenStudy (jamesj):

The domain is x such that \[ \ln x - 1 \geq 0 \] i.e., \( x \geq e \). Hence the range is \( [0, \infty) \) because the function is everywhere increasing, unbounded and evaluated at \( x = e \) the function is zero.

OpenStudy (anonymous):

How can an exponential function be zero

OpenStudy (jamesj):

Your inverse is wrong. Write the equation x = f(y) and solving for y, we have \[ x = \sqrt{\ln y - 1} \] \[ x^2 + 1 = \ln y \] \[ y = e^{x^2 + 1} \] This is the inverse function

OpenStudy (anonymous):

oh ok I get it now :) that is why i was messed up

OpenStudy (anonymous):

Everyone of my problems there is such a simple answer for :(

OpenStudy (asnaseer):

Don't worry @BlingBlong - we can only learn through mistakes.

OpenStudy (jamesj):

Notice that the domain of the inverse function is the range of the function hence the range is the the subset of the reals, \( [0, \infty) \) despite the fact that the formula for the inverse makes sense for all real numbers.

OpenStudy (jamesj):

I.e., with \( f(x) = \sqrt{ \ln x - 1 } \), \[ f : [e,\infty) \rightarrow [0,\infty) \] and \[ f^{-1} : [0,\infty) \rightarrow [e,\infty) \]

OpenStudy (anonymous):

so x in the inverse function is just y in the function

OpenStudy (anonymous):

ok I got it thanks for clarifying that for me james

OpenStudy (jamesj):

right ... the way you can find the inverse of a function f is to solve the equation x = f(y) for y.